# Aerial Vision-and-Language Navigation via Semantic-Topo-Metric Representation Guided LLM Reasoning

### v1

@@ Abstract
Aerial Vision-and-Language Navigation (VLN) is a novel task enabling Unmanned Aerial Vehicles (UAVs) to navigate in outdoor environments through natural language instructions and visual cues. It remains challenging due to the complex spatial relationships in outdoor aerial scenes. In this paper, we propose an end-to-end zero-shot framework for aerial VLN tasks, where the large language model (LLM) is introduced as our agent for action prediction. Specifically, we develop a novel Semantic-Topo-Metric Representation (STMR) to enhance the spatial reasoning ability of LLMs. This is achieved by extracting and projecting instruction-related semantic masks of landmarks into a top-down map that contains the location information of surrounding landmarks. Further, this map is transformed into a matrix representation with distance metrics as the text prompt to the LLM, for action prediction according to the instruction. Experiments conducted in real and simulation environments have successfully proved the effectiveness and robustness of our method, achieving 15.9% and 12.5% improvements (absolute) in Oracle Success Rate (OSR) on AerialVLN-S dataset.
@@ I Introduction
Figure 1: The pipeline to obtain STMR. Observation image, semantic mask, and depth image (a) will be aggregated in the top-down map (red dashed box) (b), which captures the Semantic-Topo-Metric information of the neighborhood of the current position, further generating a 20x20 matrix representation (c).
@@
To overcome these challenges, we propose a zero-shot LLM-based aerial VLN framework, which takes natural language instructions, RGB images, and depth images as input, and directly makes action predictions (e.g., forward 10 meters) through single-step reasoning updates. Specifically, a Semantic-Topo-Metric Representation (STMR) is designed for LLMs prompting. We first extract instruction-related landmarks and obtain corresponding semantic masks by perception models, i.e., Grounding Dino [liu2023grounding] and Tokenize Anything [pan2023tap]. After that, semantic masks obtained from each step are gradually projected into a top-down map using depth and pose transformation as shown in Fig. 1 (b). This projected map encompasses both the UAV’s travel trajectory and spatial information, i.e., locations of surrounding objects. To integrate rich visual information and topology into the text prompt while maintaining its simplicity, we separate the top-down map centered at the UAV’s current position into grids and substitute each grid with a semantic number. This process transforms the top-down map into a matrix representation, tailored to serve as a spatial prompt for the LLM. As can be seen from Fig. 1 (c), the matrix representation encompasses topological, semantic, and metric information, and is input to an LLM together with historical actions, sub-goal instructions, and other information to infer the next action. Compared with other learning-based and LLM-based methods, the proposed method effectively improves the accuracy of spatial reasoning.
@@
  • To our knowledge, we propose the first LLM-based end-to-end framework for the aerial VLN task, facilitating the development of UAV navigators. This framework requires no training and no extra low-level action planner, thus allowing for easy integration.
  • We propose the Semantic-Topo-Metric Representation (STMR), a unique matrix representation that encompasses topological, semantic, and metric information. STMR is designed to significantly enhance the spatial-aware reasoning capabilities of LLMs in outdoor environments.
  • Extensive experiments on the aerial VLN task especially in real environments demonstrate that the proposed method outperforms previous state-of-the-art methods by a large margin, establishing a strong baseline for future zero-shot Aerial VLN tasks.
@@ II Related Work
VLN aims to enable autonomous agents to navigate complex environments by understanding and executing natural language instructions based on visual context. Early VLN methods use sequence-to-sequence LSTMs to predict low-level actions [anderson2018vision] or high-level actions from panoramas [fried2018speaker]. Several attention processes have been proposed [qi2020object, hong2020language, an2021neighbor] to enhance the process of learning visual textual correspondence. Reinforcement learning is also explored to improve policy learning [wang2018look, tan2019learning, wang2020soft]. Besides, transformer-based architecture have shown superior performance to long-distance contextual information [hao2020towards, majumdar2020improving]. More recent works [zhou2024navgpt, chen2024mapgpt] leverage the reasoning and dialogue capabilities of LLMs, achieving great progress. However, most of them are developed under a ground-based, discrete VLN setting, which simplifies navigation as traversing on a predefined graph on the ground. This limits the free movement space of UAVs in the real world. Aerial VLN remains challenging due to the large-scale and complex environments.
@@
Unmanned Aerial Vehicle (UAV) navigation has seen a surge of interest over the past few decades. Guided with instructions, Blukis et al. [blukis2018following] use imitation learning for low-level velocity command prediction. [blukis2018mapping] predicts interpretable position-visitation distributions to guide UAV’s control actions, using a combination of supervised and imitation learning for efficient training. Misra et al. [misra2018mapping] decompose instruction execution in two stages, separately using supervised learning for goal prediction and policy gradient for action generation. AerialVLN [liu_2023_AerialVLN] contributes a much more challenging aerial VLN dataset focusing on large-scale environments, and provides a look-ahead guidance method as the baseline. Despite the progress, the generalizability and performance of these methods still require improvement.
@@ III-A Problem Formulation
In an aerial VLN task, the problem is formulated as a free-form language instruction guided navigation. At the beginning of each episode, the UAV is placed in an initial pose $P=[x,y,z,\phi,\theta,\psi]$, where $(x,y,z)$ denotes the UAV’s position and $(\phi,\theta,\psi)$ represents pitch, roll, and yaw of the UAV’s orientation. A natural language instruction $L$ is provided to specify the path that a UAV should follow. To achieve the navigation goal, the UAV considers both the instruction and visual perceptions, and predicts an action from the action space (right, left, up, down, forward, backward) with corresponding parameters at each time step $t$. Navigation ends when the UAV predicts a ’Stop’ action or reaches a pre-defined maximum action number. The navigation is recognized as a success if the UAV stops at a location that is less than 20 meters from the target location.
Figure 2: Our method consists of three modules, i.e. sub-goal extracting, Semantic-Topo-Metric Representating, and LLM planner. First, we prompt an LLM to break down long instructions and extract specific landmarks, which will be classified by perception models, generating corresponding semantic masks. Then we project these masks into a top-down map using the UAV-view pose and depth information. The map is further compressed into a matrix representation. Finally, we prompt LLM with well-designed instructions to enable effective reasoning.
@@ III-B Semantic-Topo-Metric Representation (STMR)
Previous LLM-based VLN methods use natural language to describe current observations, or nodes and edges of a topological graph to model the spatial information of the environment. However, in open scenarios, simple directional words such as "next to" or "aside" are not enough to describe complex spatial relationships, which can easily introduce ambiguity into LLMs. To address this challenge, we introduce the Semantic-Topo-Metric Representation (STMR) to enhance the spatial-aware reasoning capability of LLMs. Specifically, STMR incrementally takes an RGB image $I_{t}^{R}$ and a depth map $I_{t}^{D}$ as input from each step, and generates a dynamically updated matrix representation with semantic, topological, and metric information as its output. The details of STMR are presented as follows.
2D Visual Perception
Impressed by the powerful open-vocabulary detection capabilities of Grounding DINO $\operatorname{GD}(\cdot)$, as well as the captioning and segmentation capabilities of Tokenize Anything model $\operatorname{TA}(\cdot)$, we integrate these two models as our 2D visual perceptor, as illustrated in Fig. 3. Given a single RGB image $I_{t}^{R}$ and an instruction $L$ as input, we first obtain detailed landmark categories using a Landmark Extractor $\operatorname{LE}(\cdot)$ driven by an LLM, and then identify the corresponding bounding box for each category through $\operatorname{GD}(\cdot)$. Next, we employ $\operatorname{TA}(\cdot)$ to take each bounding box as a prompt and output a set of 2D semantic masks $m^{(t)}$ and captions $z^{(t)}$ for the current RGB image $I_{t}^{R}$. The entire process can be described as:
@@
$\{m^{(t)},\ z^{(t)}\}=\operatorname{TA}\left({I_{t}^{R}},\ \operatorname{GD}\left({I_{t}^{R}},\ \operatorname{LE}({L})\right)\right).$ (1)
To improve the simplicity of the semantic masks and reduce the misleading of LLM-reasoning caused by numerous open-vocabulary categories, we propose a text-matching method to mitigate over-classification. As shown in Fig. 3, our method involves vectorizing the landmarks extracted from the instruction and the captions $z^{(t)}$ generated in each $I_{t}^{R}$. Then, we calculate the cosine similarity between these vectorized landmarks and captions using TF-IDF [qaiser2018text]. If the similarity score exceeds the threshold $\tau$ >0.8, the landmark is classified as visible in the current view. By implementing this strategy, we effectively simplify semantic masks and ensure the LLM’s reasoning focuses on relevant categories.
Sub-goal-driven top-down map
Considering that the top-down view can better express spatial relationships than the first-person view, we further utilize the depth image $I_{t}^{D}$ to map semantic masks to 3D space, and then convert them into a top-down map. Specifically, with the RGB image $I_{t}^{R}$ processed by the 2D visual perceptor, we produce first-person view predictions by assigning semantic labels to each pixel, resulting in a segmented image with identified objects and regions. Subsequently, the depth image $I_{t}^{D}$ is converted into a 3D point cloud, where each pixel is mapped to a 3D point $(X,Y,Z)$ based on its depth value and camera parameters:
$X=\frac{(u-c_{x})I_{t}^{D}(u,v)}{f_{x}},Y=\frac{(v-c_{y})I_{t}^{D}(u,v)}{f_{y}},Z=I_{t}^{D}(u,v),$ (2)
@@
Matrix representation
We found that directly inputting the image-format top-down map to a vision-language model (VLM) often yields poor reasoning results. Therefore, we process the visual map into a text-based matrix representation, which is input to an LLM to guide the navigation. Specifically, the UAV incrementally collects observations and gradually constructs the sub-goal-driven top-down map of the surrounding environment. Considering the perception quality of onboard sensors and the effective processing capacity of modern LLMs, we define a local map centered at the UAV’s current position as shown in Fig. 1. Subsequently, we generate a matrix by partitioning the local map into a $20\times 20$ coordinate grid and applying semantic max pooling in each grid, making them finally contain one semantic category. Particularly, we grade all values that fall in each grid in the local map according to frequency, showing only the most frequent features as the final category of this grid. Each category such as buildings or cars are discriminated by different number labels. Furthermore, the distance between each grid, denoted as $r$ (5m), is also input to the LLM as the metric information. In this case, the LLM can directly reason about positional relationships among each landmark.
@@ III-C Designing Text Prompt for LLM Planning
Given that LLMs are well-trained in reasoning and planning, we complement their spatial perception and scene understanding capability with formatted prompt and action space. This prompt involves two components. First, the task description is prepared which includes the definition of the environment and the UAV, as well as the format of input and output. We require an LLM to leverage Chain-of-Thought reasoning [wei2022chain] after capturing the observation, thinking step by step in an order of observation-thought-planning-history to predict the next action. Regarding the input instruction, sub-goals are extracted by the LLM, and the landmarks given in the instructions will be updated in the task definition as the scene changes. Second, to improve the quality of prompts, we implemented a cyclic refinement process. This involves designing a new set of instructions for an LLM, including elements such as ‘history action’ and ‘next planning’ corresponding to the STMR module. The task description also emphasizes updating the status of planning, including three states, i.e., todo, in process, and completed. The STMR is gradually updated in this iterative process as the UAV explores to improve the accuracy and relevance of the prompts.
Example of LLM Prompt [Task Description]You are an embodied unmanned aerial vehicle that navigates in the real world. You need to explore between some places marked and ultimately find the destination to stop.
[Input Format]: ‘Instruction’ is a global, step-by-step detailed guidance. ‘History’ is your previously executed actions. ‘Map’ is a top-down coordinate map, using a 20*20 matrix M to represent coordinates. The numbers in the matrix represent different landmarks. You are at [10,10] of the matrix, which is a special value denoting your orientation, e.g., west0 means face due west and the declination angle is 0 degree. The distance between adjacent numbers is 5 meters in the real world. The numbers in the matrix M represent:
[0:Unexplored 1:road 2:building … -1:your past trajectory]. ‘Plan’ records previous long-term multi-step planning info that you can refer to now.
[Output Format] ‘Thought’: your thoughts about the task, which may include your comprehension, surroundings, history, and etc. ‘Observation’: the description of the map. ‘Plan’: your updated plan. ‘Action’: your next action."
Think step by step. First, judge by the ‘Map’, give a first ‘Thought’, and depict your orientation. Second, check that if a landmark in the current ‘plan’ is within 5 meters of your current position, then based on ‘Instruction’ and the previous ‘Plan’, update your multi-step ‘Plan’. Each plan needs to follow a state word (TODO, In Process, Completed). Finally, judge by the ‘Map’ again, and give the specific "Action" according to the action format. Action format: Action: (right, left, lift, down, straight, back), (degree (0-15 degrees)), (distance (0-10 meters)).
  Instruction: lift off and head straight across the water to the road…

### v2

# Exploring Spatial Representation to Enhance LLM Reasoning in Aerial Vision-Language Navigation

@@ Abstract
Aerial Vision-and-Language Navigation (VLN) is a novel task enabling Unmanned Aerial Vehicles (UAVs) to navigate in outdoor environments through natural language instructions and visual cues. However, it remains challenging due to the complex spatial relationships in outdoor aerial scenes. In this paper, we propose a training-free, zero-shot framework for aerial VLN tasks, where the large language model (LLM) is leveraged as the agent for action prediction. Specifically, we develop a novel Semantic-Topo-Metric Representation (STMR) to enhance the spatial reasoning capabilities of LLMs. This is achieved by extracting and projecting instruction-related semantic masks of landmarks onto a top-down map, which contains location information about surrounding landmarks. The map is then transformed into a matrix representation with distance metrics, serving as the text prompt to the LLM for action prediction in response to the given instruction. Experiments conducted in real and simulation environments have successfully proved the effectiveness and robustness of our method, achieving at most 10.8% improvements (absolute) in Oracle Success Rate (OSR) on AerialVLN-S and Openfly dataset.
@@ I Introduction
Figure 1: The pipeline to obtain STMR. Observation image, semantic mask, and depth image (a) can be aggregated in the top-down map (red dashed box) (b), which captures the Semantic-Topo-Metric information of the neighborhood of the current position, further generating a 20x20 matrix representation (c) for LLM reasoning.
@@
To overcome these challenges, we propose a zero-shot LLM-based aerial VLN framework, which takes natural language instructions, RGB images, and depth images as input, and directly makes action predictions (e.g., forward 10 meters) through single-step reasoning updates. Specifically, a Semantic-Topo-Metric Representation (STMR) is designed for LLMs spatial reasoning. Leveraging the developments in open-vocabulary perception models [zhang2018review, jiang2018deep, minderer2022simple, cheng2024yolo], we first extract instruction-related landmarks and obtain corresponding semantic masks by perception models, i.e., Grounding Dino [liu2023grounding] and Tokenize Anything [pan2023tap]. After that, semantic masks obtained from each step are gradually projected into a top-down map using depth and pose transformation as shown in Fig. 1 (b). This projected map encompasses both the UAV’s travel trajectory and spatial information, i.e., locations of surrounding objects. To integrate rich visual information and topology into the text prompt while maintaining its simplicity, we separate the top-down map centered at the UAV’s current position into grids and substitute each grid with a semantic number. This process transforms the top-down map into a matrix representation, tailored to serve as a spatial prompt for the LLM. As can be seen from Fig. 1 (c), the matrix representation encompasses topological, semantic, and metric information, and is input to an LLM together with historical actions, sub-goal instructions, and other information to infer the next action. Compared with other learning-based and LLM-based methods, the proposed method effectively improves the accuracy of spatial reasoning.
@@
  • To our knowledge, we design the first LLM-based training-free framework for the aerial VLN task, facilitating the development of UAV navigators. Without extra action planner, the proposed framework allows for easy integration.
  • We propose the Semantic-Topo-Metric Representation (STMR), a unique matrix representation that encompasses topological, semantic, and metric information. STMR is designed to enhance the spatial-aware reasoning capabilities of LLMs in outdoor environments.
  • Extensive experiments on the aerial VLN task demonstrate that the proposed method outperforms previous state-of-the-art methods by a large margin, establishing a strong baseline for future zero-shot Aerial VLN tasks.
@@ II Related Work
VLN aims to enable autonomous agents to navigate complex environments by understanding and executing natural language instructions based on visual context. Early VLN methods use sequence-to-sequence LSTMs to predict low-level actions [anderson2018vision] or high-level actions from panoramas [fried2018speaker]. Several attention processes have been proposed [qi2020object, hong2020language, an2021neighbor] to enhance the process of learning visual textual correspondence. Reinforcement learning is also explored to improve policy learning [wang2018look, tan2019learning, wang2020soft]. Besides, transformer-based architecture have shown superior performance to long-distance contextual information [hao2020towards, majumdar2020improving]. More recent works [zhou2024navgpt, chen2024mapgpt] leverage the reasoning and dialogue capabilities of LLMs, achieving great progress. However, most of them are developed under a ground-based, discrete VLN setting, which simplifies navigation by traversing on a predefined graph on the ground. This limits the UAVs’ free movement space in the real world. Aerial VLN remains challenging due to the large-scale and complex environments.
@@
Unmanned Aerial Vehicle (UAV) navigation has seen a surge of interest over the past few decades. Guided with instructions, Blukis et al. [blukis2018following] use imitation learning for low-level velocity command prediction. [blukis2018mapping] predicts interpretable position-visitation distributions to guide UAV’s control actions, using a combination of supervised and imitation learning for efficient training. Misra et al. [misra2018mapping] decompose instruction execution in two stages, separately using supervised learning for goal prediction and policy gradient for action generation. AerialVLN [liu_2023_AerialVLN] contributes a much more challenging aerial VLN dataset focusing on outdoor aerial environments, and provides a look-ahead guidance method as the baseline. OpenFly [openfly] constructs a large-scale dataset across several high-quality scenes, and proposes a keyframe-aware VLN model. Despite the progress, the generalizability and performance of these methods still require improvement.
@@ III-A Problem Formulation
In an aerial VLN task, the problem is formulated as a free-form language instruction guided navigation. At the beginning of each episode, the UAV is placed in an initial pose $P=[x,y,z,\phi,\theta,\psi]$, where $(x,y,z)$ denotes the UAV’s position and $(\phi,\theta,\psi)$ represents pitch, roll, and yaw of the UAV’s orientation. A natural language instruction $L$ is provided to specify the path that a UAV should follow. To achieve the navigation goal, the UAV considers both the instruction and visual perceptions, and predicts an action from the action space (right, left, up, down, forward, backward) with corresponding parameters at each time step $t$. Navigation ends when the UAV predicts a ‘Stop’ action or reaches a pre-defined maximum action number. Following [liu_2023_AerialVLN], the navigation is recognized as a success if the UAV stops at a location that is less than 20 meters from the target location.
Figure 2: Our method consists of three modules, i.e. sub-goal extracting, Semantic-Topo-Metric Representation, and LLM planner. First, we prompt an LLM to break down long instructions and extract specific landmarks, which will be classified by perception models, generating corresponding semantic masks. Then we project these masks into a top-down map using the UAV-view pose and depth information. The map is further compressed into a matrix representation. Finally, we prompt LLM with well-designed instructions to enable effective reasoning.
@@ III-B Semantic-Topo-Metric Representation (STMR)
Previous LLM-based VLN methods use natural language to describe current observations, or nodes and edges of a topological graph to model the spatial information of the environment. However, in open scenarios, simple directional words such as “next to” or “aside” are not enough to describe complex spatial relationships, which can easily introduce ambiguity into LLMs. To address this challenge, we introduce the STMR to enhance the spatial-aware reasoning capability of LLMs. Specifically, STMR incrementally takes an RGB image $I_{t}^{R}$ and a depth map $I_{t}^{D}$ as input from each step, and generates a dynamically updated matrix representation with semantic, topological, and metric information as its output. The details of STMR are presented as follows.
III-B1 2D Visual Perception
Impressed by the powerful open-vocabulary detection capabilities of Grounding DINO $\operatorname{GD}(\cdot)$, as well as the captioning and segmentation capabilities of Tokenize Anything model $\operatorname{TA}(\cdot)$, we integrate these two models as our 2D visual preceptor, as illustrated in Fig. 3. Given a single RGB image $I_{t}^{R}$ and an instruction $L$ as input, we first obtain detailed landmark categories $C=\{c_{1},c_{2},...,c_{n}\}$ using a Landmark Extractor $\operatorname{LE}(\cdot)$ driven by an LLM, and then identify the corresponding bounding box for each category through $\operatorname{GD}(\cdot)$. Next, we employ $\operatorname{TA}(\cdot)$ to take each bounding box as a prompt and output a set of 2D semantic masks $m^{(t)}$ and captions $h^{(t)}$ for the current RGB image $I_{t}^{R}$. The entire process can be described as:
@@
$\{m^{(t)},\ h^{(t)}\}=\operatorname{TA}\left({I_{t}^{R}},\ \operatorname{GD}\left({I_{t}^{R}},\ \operatorname{LE}({L})\right)\right).\$ (1)
To improve the simplicity of the semantic masks and reduce the misleading of LLM reasoning caused by numerous open-vocabulary categories, we propose a text-matching method to mitigate over-classification. As shown in Fig. 3, our method involves vectorizing the landmarks extracted from the instruction and the captions $h^{(t)}$ generated in each $I_{t}^{R}$. Then, we calculate the cosine similarity between these vectorized landmarks and captions using TF-IDF [qaiser2018text]. If the similarity score exceeds the threshold $\tau$ (0.8), the landmark is classified as visible in the current view, and other irrelevant masks will be ignored. By implementing this strategy, we effectively simplify semantic masks and ensure the LLM’s reasoning focuses on relevant categories.
III-B2 Sub-goal-driven top-down map
Considering that the top-down view can better express spatial relationships than the first-person view, we further utilize the depth image $I_{t}^{D}$ to map semantic masks to 3D space, and then convert them into a top-down map. Specifically, by applying the above 2D visual perception process to the RGB image $I_{t}^{R}$, we obtain a segmented image with clearly identified objects and regions. Subsequently, the depth image $I_{t}^{D}$ is converted into a 3D point cloud, where each pixel is mapped to a corresponding 3D point $(X,Y,Z)$ based on its depth value and camera parameters:
$X=\frac{(u-c_{x})I_{t}^{D}(u,v)}{f_{x}},Y=\frac{(v-c_{y})I_{t}^{D}(u,v)}{f_{y}},$ (2)
$Z=I_{t}^{D}(u,v),$ (3)
@@ III-B2 Sub-goal-driven top-down map
III-B3 Matrix representation
We found that directly inputting the image-format top-down map to a vision-language model (VLM) often yields poor reasoning results. Alternatively, we process the visual map into a text-based matrix representation, which is input to an LLM to guide the navigation. Specifically, the UAV incrementally collects observations and gradually constructs the top-down map of the surrounding environment. Considering the perception quality of onboard sensors and the effective processing capacity of modern LLMs, we define a local map centered at the UAV’s current position as shown in Fig. 1. Subsequently, we generate a matrix by partitioning the local map into a $20\times 20$ coordinate grid and applying semantic max pooling in each grid, making them finally contain one semantic category. Particularly, we grade all values that fall in each grid in the local map according to frequency, showing only the most frequent features as the final category of this grid. Each category such as building or car is discriminated by different number labels. Furthermore, the grid size $r$ (5m) is also input to the LLM as the metric information (see Table VII for ablation study on the grid size). In this case, the LLM can directly reason about positional relationships among each landmark.
@@ III-C Designing Text Prompt for LLM Planning
Given that LLMs are well-trained in reasoning and planning, we complement their spatial perception and scene understanding capability with formatted prompt and action space. This prompt involves two components. First, the task description is prepared which includes the definition of the environment and the UAV, as well as the format of input and output. We require an LLM to leverage Chain-of-Thought reasoning [wei2022chain] after capturing the observation, thinking step by step in an order of observation-thought-planning-history to predict the next action. Regarding the input instruction, sub-goals are extracted by the LLM, and the landmarks given in the instructions will be updated in the task definition as the scene changes. Second, to improve the quality of prompts, we implemented a cyclic refinement process. This involves designing a new set of instructions for an LLM, including elements such as ‘history action’ and ‘next planning’ corresponding to the STMR module. The task description also emphasizes updating the status of planning, including three states, i.e., todo, in process, and completed, in order to minimize the problem of hallucinations that is common in LLM. The STMR is gradually updated in this iterative process as the UAV explores to improve the accuracy and relevance of the prompts.
Example of LLM Prompt (simplified) [Task Description]You are an embodied unmanned aerial vehicle that navigates in the real world. You need to explore between some places marked and ultimately find the destination to stop.
[Input Format] ‘Instruction’ is a global, step-by-step detailed guidance. ‘History’ is your previously executed actions. ‘Map’ is a top-down coordinate map, using a 20*20 grid M to represent coordinates, heading towards upper of the map. The value of the grid represent different landmarks. You are at [9,9] on the ’Map’ (marked as P). The distance between adjacent numbers is 5 meters in the real world. The numbers in the matrix M represent:
[0:Unexplored 1:your past trajectory 2:road 3:building … ]. ‘Plan’ records previous long-term multi-step planning info that you can refer to now.
[Output Format] ‘Thought’: your thoughts about the task, which may include your comprehension, surroundings, history, and etc. ‘Observation’: the description of the map. ‘Plan’: your updated plan. ‘Action’: your next action.”
Think step by step. First, describe your orientation and the landmark in your left, right and front. Second, according the ’map’, update your multi-step ‘Plan’. Each plan needs to follow a state word (TODO, In Process, Completed). Finally, give the specific “Action” according to the action format: Action: (right, left, lift, down, straight, back), (degree (0-90 degrees)), (distance (between 0-10 meters)).
  A Simplified Input in the Cyclic Process
Instruction
: lift off and head straight across the water to the road…